# A Search for Low-Mass Neutron Stars in the Third Observing Run of Advanced LIGO and Virgo

### v1

@@ Abstract
Most observed neutron stars have masses around 1.4 Mdirect-$\odot$, consistent with current formation mechanisms. To date, no sub-solar mass neutron star has been observed. Observing a low-mass neutron star would be a significant milestone, providing crucial constraints on the nuclear equation of state, unveiling a new population of neutron stars, and advancing the study of their formation processes and underlying mechanisms. We present the first targeted search for tidally deformed sub-solar mass binary neutron stars (BNS), with primary masses ranging from 0.1 to 2 Mdirect-$\odot$ and secondary masses from 0.1 to 1 Mdirect-$\odot$, using data from the third observing run of the Advanced LIGO and Advanced Virgo gravitational-wave detectors. We account for the tidal deformabilities of up to $O(10^{4})$ of these systems, as low-mass neutron stars are more easily distorted by their companions’ gravitational forces. Previous searches that neglect tidal deformability lose sensitivity to low-mass sources, potentially missing more than $\sim 30\%$ of detectable signals from a system with a chirp mass of 0.6 Mdirect-$\odot$ binaries. No statistically significant detections were made. In the absence of a detection, we place a 90% confidence upper limit on the local merger rate for sub-solar mass BNS systems, constraining it to be $<6.4\times 10^{4}$ Gpc-3Yr-1 for a chirp mass of 0.2 Mdirect-$\odot$ and $<2.2\times 10^{3}$ Gpc-3Yr-1 for 0.7 Mdirect-$\odot$. With future upgrades to detector sensitivity, development of next-generation detectors, and ongoing improvements in search pipelines, constraints on the minimum mass of neutron stars will improve, providing the potential to constrain the nuclear equation of state, reveal new insights into neutron star formation channels, and potentially identify new classes of stars.
@@ 1 Introduction
One active area of research focuses on constraining the nuclear equation of state (EOS) of neutron stars. EOS models suggest that a neutron star with a mass of 1 Mdirect-$\odot$ could have a radius between 11 and 15 km (Abbott et al., 2020a; Baldo et al., 1997; Müller & Serot, 1996; Danielewicz & Lee, 2009; Agrawal et al., 2005; Nazarewicz et al., 1996), a 2 Mdirect-$\odot$ neutron star may have a radius from 10 to 14 km, and at the extreme ends a 0.4Mdirect-$\odot$ neutron star may have radii anywhere from 11 km to 16.5 km. The wide range reflects the significant uncertainty in EOS constraints at the extreme densities, resulting in multiple competing models (Biswas & Rosswog, 2024; Chatziioannou et al., 2024; Özel & Freire, 2016; Baym et al., 2018). At the upper mass limit, the mass-radius curves of the EOS show that the neutron star is predicted to have a smaller radius than a low-mass neutron star. Extending the EOS to higher masses is difficult due to causality constraints, as the pressure-density relationship would require the speed of sound to exceed the speed of light for masses above 2 Mdirect-$\odot$ (Sumiyoshi et al., 2022). At the lower mass limit, some EOS models describe neutron stars as light as 0.1 Mdirect-$\odot$, though few plausible formation mechanisms exist for these stars, and no direct observations have yet confirmed their existence (Lattimer, 2021; The LIGO Scientific Collaboration et al., 2022; Nitz & Wang, 2022; Metzger et al., 2024). Overall, constraints on the EOS, particularly at the extreme mass limits, remain reliant on observations and are likely to improve as more data becomes available.
The first observation of a neutron star began with the detection of a radio pulsar in 1967. Since then, more than 2,000 neutron stars have been cataloged (Lattimer & Prakash, 2005). The MPIfR-MeerKAT Galactic Plane Survey cataloged 78 radio pulsars or rapidly rotating, highly magnetized systems, 17 of which were identified as BNSs (Colom i Bernadich et al., 2023), while the Fermi Large Area Telescope has identified approximately 300 gamma-ray pulsars (Smith et al., 2023). X-ray observations from NICER have expanded the neutron star population to include around 947 X-ray pulsars (NASA High Energy Astrophysics Science Archive Research Center (2024), HEASARC). Additionally, the Konus-Wind spectrometer has observed 199 short gamma-ray bursts, which result from violent mergers of compact objects (Lysenko et al., 2024). Most of the observed neutron stars have masses near 1.4 Mdirect-$\odot$ (Lattimer & Prakash, 2005). In addition to electromagnetic observations, there have been detections of BNS using gravitational waves (GWs) (Abbott et al., 2017a, 2020b).
@@
GWs are distortions in spacetime produced by the acceleration of massive objects. These waves can be generated from binary systems of compact objects, such as neutron stars and black holes, where GWs are radiated as the two compact objects inspiral toward each other and eventually merge (Dirkes, 2018). Ground-based observatories such as Advanced LIGO and Advanced Virgo measure the tiny distortions caused on the detector by the GWs of the compact binary (Acernese et al., 2015; LIGO Scientific Collaboration et al., 2015). To date, two BNS mergers have been detected via GWs (Abbott et al., 2017a, 2020b). The first detection, GW170817, had a chirp mass of 1.186 Mdirect-$\odot$ and was accompanied by an electromagnetic counterpart (Abbott et al., 2017b; Alexander et al., 2017; Abbott et al., 2017; Valenti et al., 2017). This event confirmed that BNS collisions are the progenitors of kilonovae and gamma-ray bursts (Abbott et al., 2017a, 2017).
@@
GW observations allow for the measurement of the masses and tidal distortions of neutron stars through their imprint on the gravitational waveform. The masses influence the amplitude and frequency of the signal, with larger masses increasing the amplitude of the waveform and shifting the frequency to lower values. Additionally, the phase evolution of the waveform encodes information about the tidal distortions of the stars, often referred to as tidal deformability. Together, the masses and tidal deformability provide direct constraints on the EOS, making GW observations a powerful tool for probing the EOS of neutron stars. Tidal deformability $\lambda$ is defined as
@@
If the detected object is indeed a neutron star, its low mass suggests it likely formed through alternative mechanisms, as traditional supernovae are not expected to produce such low-mass remnants (no less than 1.2 Mdirect-$\odot$) (Kiziltan et al., 2013; Lattimer & Prakash, 2001; Douchin & Haensel, 2001; Lattimer, 2012; Suwa et al., 2018). One proposed mechanism for generating a low-mass neutron star involves the fragmentation of the accretion disk around a black hole (Metzger et al., 2024). In this scenario, the collapse of a rapidly rotating massive star (greater than 20 Mdirect-$\odot$) leads to the formation of a centrifugally supported torus. Matter from the outer layers of the star’s core fragments into multiple pieces, which can then condense into sub-solar mass neutron stars. These systems, which are expected to be in binary configurations, could be detected by ground-based gravitational wave detectors such as Advanced LIGO and Advanced Virgo.
@@ 2 Search
Stochastic template bank methods do not guarantee recovery of all possible signals at or above the minimal match threshold, potentially leaving gaps in coverage. We validated the bank’s completeness by randomly generating waveform models and calculating matches between injected signals and their nearest templates. The bank was designed such that 99% of templates successfully recover SNRs $\geq 95\%$ for fiducial injections within the parameter space. However, coverage was weakest for equal-mass systems at 0.1 Mdirect-$\odot$. Less than 1% of the bank failed to recover matches $\geq 0.95$. The final bank contained approximately $10^{7}$ templates.
@@ 3 Results
From our missed detection criteria we can calculate the upper limit of the merger rate $\mathcal{R}(\theta,z)$ for a given astrophysical model. To constrain the upper limit of the merger rate given our set of populations, the expected number of detected events $N$ is given by
@@
Where $d\theta+dz=dV$, $\mathcal{R}(\theta,z)$ is the density merger rate, $dt$ is the period of observing time, $\epsilon(\theta,z)$ is the detection efficiency which is the probability of measuring a detection within a given set of parameters $\theta$, and redshift z. Since our search is only sensitive to nearby sources, we do not consider the cosmological effects. The number of detected events will be reduced to
@@
2.303 is the expected number of detections given a 90% confidence drawn from a Poisson distribution (Biswas et al., 2009).
@@
We calculated the averaged volume time product $\langle VT\rangle$ using a Monte Carlo integral method for the injected population of signals by counting the number of missed and found injections from our search pipeline. We injected the signals assuming an isotropic distribution, with sources randomly distributed across sky positions and orientations. We fixed five chirp masses and five tidal deformabilities. Figure 3 shows the upper limit for these fixed parameters. Our search is equally sensitive across all tidal deformabilities. Overall, we find that our search constrains the upper limit of the merger rate for binary neutron stars to be $<6.4\times 10^{4}$ Gpc-3Yr-1 for chirp mass 0.2 Mdirect-$\odot$ and $<2.2\times 10^{3}$ Gpc-3Yr-1 for 0.7 Mdirect-$\odot$.
4 Conclusion
@@
This discovery could also drive the investigation of exotic matter states, such as quark matter and dark matter, potentially revealing unique material properties under extreme densities and pressures. If an exotic star, such as a quark star, can be distinguished from a neutron star, it may offer new opportunities for detecting dark matter signatures (Hong & Ren, 2024). Previous research suggests that a quark star’s strong gravitational fields might capture dark matter (Hong & Ren, 2024). Observing GWs from low-mass stars that interact with dark matter could offer a novel way to measure dark matter directly, as such stars might contain dark matter components that influence their structure and tidal deformability.
This search provided several insights for computational challenges for future low-mass searches. One of the main limiting factors is the number of templates used. Our bank covered the potential tidal deformable range of up to $\mathcal{O}(10^{4})$ for a 0.1 Mdirect-$\odot$ neutron star. Extending the analysis to cover the full tidal deformability range (up to $\mathcal{O}(10^{6})$) was not feasible due to the rapid growth of the template bank. We also limited our search to the most sensitive data publicly available. Future searches may revisit this search by incorporating the full tidal deformability range and performing the search over all the current and future available data. To accomplish this in a reasonable amount of time, future searches will require new methodologies to handle larger, more complex parameter spaces efficiently.

### v2

@@ Abstract
Most observed neutron stars have masses around 1.4 Mdirect-$\odot$, consistent with current formation mechanisms. To date, no sub-solar mass neutron star has been observed. Observing one would provide crucial constraints on the nuclear equation of state, unveil a new neutron star population, and advance understanding of their formation mechanisms. We present the first targeted search for tidally deformed sub-solar mass binary neutron stars (BNS), with primary masses ranging from 0.1 to 2 Mdirect-$\odot$ and secondary masses from 0.1 to 1 Mdirect-$\odot$, using data from the third observing run of the Advanced LIGO and Virgo gravitational-wave detectors. We account for the tidal deformabilities of up to $O(10^{4})$ of these systems, as low-mass neutron stars are more easily distorted by their companion. Previous searches that neglect tidal deformability lose sensitivity to low-mass sources, potentially missing more than $\sim 30\%$ of detectable signals from a system with a chirp mass of 0.6 Mdirect-$\odot$ binaries. No statistically significant detections were made. In the absence of a detection, we place a 90% confidence upper limit on the local merger rate for sub-solar BNSs, constraining it to be $<6.4\times 10^{4}$ Gpc-3Yr-1 for a chirp mass of 0.2 Mdirect-$\odot$ and $<2.2\times 10^{3}$ Gpc-3Yr-1 for 0.7 Mdirect-$\odot$. With future upgrades to detector sensitivity, development of next-generation detectors, and ongoing improvements in search pipelines, constraints on the minimum mass of neutron stars will improve, providing the potential to constrain the nuclear equation of state, reveal new insights into neutron star formation channels, and potentially identify new classes of stars.
@@ 1 Introduction
One active area of research focuses on constraining the nuclear equation of state (EOS) of neutron stars. EOS models suggest that a neutron star with a mass of 1 Mdirect-$\odot$ could have a radius between 11 and 15 km (Abbott et al., 2020a; Baldo et al., 1997; Müller & Serot, 1996; Danielewicz & Lee, 2009; Agrawal et al., 2005; Nazarewicz et al., 1996), a 2 Mdirect-$\odot$ neutron star may have a radius from 10 to 14 km, and at the extreme ends a 0.4Mdirect-$\odot$ neutron star may have radii anywhere from 11 km to 16.5 km. The wide range reflects the significant uncertainty in EOS constraints at the extreme densities, resulting in multiple competing models (Biswas & Rosswog, 2024; Chatziioannou et al., 2024; Özel & Freire, 2016; Baym et al., 2018). At the upper mass limit, the mass-radius curves of the EOS show that the neutron star is predicted to have a smaller radius than a low-mass neutron star. Extending the EOS to higher masses is difficult due to causality constraints, as the pressure-density relationship would require the speed of sound to exceed the speed of light for masses above 2 Mdirect-$\odot$ (Sumiyoshi et al., 2022). At the lower mass limit, some EOS models describe neutron stars as light as 0.1 Mdirect-$\odot$, though few plausible formation mechanisms exist for these stars, and no direct observations have yet confirmed their existence (Lattimer, 2021; The LIGO Scientific Collaboration et al., 2022; Nitz & Wang, 2022; Metzger et al., 2024). Overall, constraints on the EOS, particularly at the extreme mass limits, still depend on observational data and are expected to improve as more BNS mergers are detected, including those observed through gravitational waves (GWs).
@@
GWs are distortions in spacetime produced by the acceleration of massive objects. These waves can be generated from binary systems of compact objects, such as neutron stars and black holes, where GWs are radiated as the two compact objects inspiral toward each other and eventually merge (Dirkes, 2018). Ground-based observatories such as Advanced LIGO and Advanced Virgo measure the tiny distortions caused on the detector by the GWs of the compact binary (Acernese et al., 2015; LIGO Scientific Collaboration et al., 2015). To date, two BNS mergers have been detected via GWs (Abbott et al., 2017a, 2020b). The first detection, GW170817, had a chirp mass $\mathcal{M}$ of 1.186 Mdirect-$\odot$ and was accompanied by an electromagnetic counterpart.(Abbott et al., 2017b; Alexander et al., 2017; Abbott et al., 2017; Valenti et al., 2017). This event confirmed that BNS collisions are the progenitors of kilonovae and gamma-ray bursts (Abbott et al., 2017a, 2017).
GW observations allow for the measurement of the masses and tidal distortions of neutron stars through their imprint on the gravitational waveform. The chirp mass $\mathcal{M}$ is defined as:
$\mathcal{M}=\frac{(m_{1}m_{2})^{\frac{3}{5}}}{(m_{1}+m_{2})^{\frac{1}{5}}}$ (1)
@@
where $m_{1}$ and $m_{2}$ are the masses of the two compact objects in the binary. The masses influences the amplitude and frequency of the signal, with larger masses increasing the amplitude of the waveform and shifting the frequency to lower values. Additionally, the phase evolution of the waveform encodes information about the tidal distortions of the stars, often referred to as tidal deformability. Together, the masses and tidal deformability provide direct constraints on the EOS, making GW observations a powerful tool for probing the EOS of neutron stars. Tidal deformability $\lambda$ is defined as
@@
If the detected object is indeed a neutron star, its low mass suggests it likely formed through alternative mechanisms, as traditional supernovae are not expected to produce such low-mass remnants (no less than 1.2 Mdirect-$\odot$) (Kiziltan et al., 2013; Lattimer & Prakash, 2001; Douchin & Haensel, 2001; Lattimer, 2012; Suwa et al., 2018; Crescimbeni et al., 2024b). One proposed mechanism for generating a low-mass neutron star involves the fragmentation of the accretion disk around a black hole (Metzger et al., 2024). In this scenario, the collapse of a rapidly rotating massive star (greater than 20 Mdirect-$\odot$) leads to the formation of a centrifugally supported torus. Matter from the outer layers of the star’s core fragments into multiple pieces, which can then condense into sub-solar mass neutron stars. These systems, which are expected to be in binary configurations, could be detected by ground-based gravitational wave detectors such as Advanced LIGO and Advanced Virgo.
@@ 2 Search
Stochastic template bank methods do not guarantee recovery of all possible signals at or above the minimal match threshold, potentially leaving gaps in coverage. We validated the bank’s completeness by randomly generating waveform models and calculating matches between injected signals and their nearest templates. The bank was designed such that 99% of templates successfully recover SNRs $\geq 95\%$ for fiducial injections within the parameter space. However, coverage was weakest for equal-mass systems at 0.1 Mdirect-$\odot$. Less than 1% of the bank failed to recover matches $\geq 0.95$. The final bank contained approximately $10^{7}$ templates. We also tested the bank with a more modern waveform model to evaluate any potential improvement in sensitivity. We compared the recovered SNRs from injections generated with IMRPhenomXP_NRTidalv2 to those from our final bank. We found that the recovered SNRs were consistent with the TaylorF2 injections, showing no improvement in sensitivity.
@@ 3 Results
From our missed detection criteria we can calculate the upper limit of the merger rate $\mathcal{R}(\theta,z)$ for a given astrophysical model. To constrain the upper limit of the merger rate given our set of populations, the expected number of detected events $N$ is given by (Biswas et al., 2009; Abbott et al., 2016)
@@
Where $t$ is the period of observing time, $(dV/dz)$ is the differential commoving volume, $d\mathcal{R}(\theta,\zeta,z)/d(\theta)$ is the differential merger rate density, $\epsilon(\theta,z)$ is the detection efficiency, which is the probability of measuring a detection within a given set of parameters $\theta$, and redshift z of the binary. Since our search is only sensitive to nearby sources, we do not consider the cosmological effects. The number of detected events will be reduced to
[…]
Note that in this case, $\epsilon$ still depends on redshift and $z$ is absorbed into $\theta$. For a given set of intrinsic parameter $(\theta_{0})$ we can integrate out the rate density
[…]
@@
Where 2.303 is the expected number of detections $N$ given a 90% confidence drawn from a Poisson distribution (Biswas et al., 2009).
@@
We calculated the averaged volume time product $\langle VT\rangle$ using a Monte Carlo integral method for the injected population of signals by counting the number of missed and found injections from our search pipeline. We injected the signals assuming an isotropic distribution, with sources randomly distributed across sky positions and orientations. We fixed five chirp masses and five tidal deformabilities. Figure 3 shows the upper limit for these fixed parameters. Our search is equally sensitive across all tidal deformabilities. Overall, we find that our search constrains the upper limit of the merger rate for binary neutron stars to be $<6.4\times 10^{4}$ Gpc-3Yr-1 for chirp mass 0.2 Mdirect-$\odot$ and $<2.2\times 10^{3}$ Gpc-3Yr-1 for 0.7 Mdirect-$\odot$. We find that our results are consistent with previous non-tidally deformed sub-solar mass searches. For a 1.1 chirp mass system, previous searches rates have calculated the upper limit to be $\sim 7\times 10^{2}$ (The LIGO Scientific Collaboration et al., 2022), where we have $<7\times 10^{2}$ Gpc-3Yr-1 for all O3 data and tidally deformed sources.
4 Discussions and Conclusion
@@
This discovery could also drive the investigation of exotic matter states, such as quark matter and dark matter, potentially revealing unique material properties under extreme densities and pressures. If an exotic star, such as a quark star, can be distinguished from a neutron star, it may offer new opportunities for detecting dark matter signatures (Hong & Ren, 2024). Previous research suggests that a quark star’s strong gravitational fields might capture dark matter (Hong & Ren, 2024). Observing GWs from low-mass stars that interact with dark matter could offer a novel way to measure dark matter directly, as such stars might contain dark matter components that influence their structure and tidal deformability. Conversely, if a sub-solar object is found to have no tidal deformability, it could be distinguished from stars as a primordial black hole, another intriguing theoretical candidate that may host clouds of dark matter, providing an alternative means of probing dark matter properties (Crescimbeni et al., 2024a).
This search provided several insights for computational challenges for future low-mass searches. One of the main limiting factors is the number of templates used. Our bank covered the potential tidal deformable range of up to $\mathcal{O}(10^{4})$ for a 0.1 Mdirect-$\odot$ neutron star. Extending the analysis to cover the full tidal deformability range (up to $\mathcal{O}(10^{6})$) was not feasible due to the rapid growth of the template bank. Our search was also restricted to the most sensitive publicly available data. A recent study suggests that data from the fourth and fifth observing runs of LIGO, Virgo and KAGRA could have sufficient sensitivity to detect sub-solar mass events, should they exist (Crescimbeni et al., 2024a, b). Future searches may revisit this search by incorporating the full tidal deformability range and performing the search over all the current and future available data. To accomplish this in a reasonable amount of time, future searches will require new methodologies to handle larger, more complex parameter spaces efficiently.